

********************************************************************
UARS WEEKLY STATUS REPORT - 14 November 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE was turned on November 11, 1997. On the next pass (TDRS contact) the
end-of night voltage dropped below 24 volts. The HALOE heater/motor bus
was turned off to reduce the load on the bus. The main HALOE power bus was
left on and the heater/motor bus was turned back on at approximately 16:00
GMT. The UARS FOT was very helpful during this process.
***********
MLS
***********
The MLS instrument has been powered off since Nov 4. MLS is scheduled to
begin north viewing operations on Nov 20.
***********
HRDI
***********
HRDI was turned on 11/13/97. We are scheduled to be on till 11/19/97.
Operation is normal.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
UARS CDHF Status Week Ending November 14, 1997
---------------------------------------------
The software that generates the UARS level 0 files and related NRT and Quicklook
products will be re-hosted to a new hardware platform on Saturday November 15,
1997. UARS CDHF users have been notified about the delay in NRT, Quicklook and
potentially the Level 0 product that will occur on this date. The time
estimate for this re-hosting activity is from 12:00GMT- 17:00GMT Nov. 15.
The CDHF has received a total of 341 CD-ROM requests as of 14 November 97.
One new CD-ROM order was received and shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
30 days PEM (3.32) Prod (1350R-565C=785L)
13 days MLS (4.9) ENG (40R-40C=0L)
The CDHF has previewed the DAAC Transfer Data for September 1997 current day
data.
Pending CDHF software upgrades include IDL 5.0, and OpenIngres 2.0
***********
FOT
***********
REPORTING PERIOD
This report covers 31 October 1997 (Orbit 33548, GMT Day 304) through
07 November 1997 (Orbit 33653, GMT Day 311).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 306/0900 33570 ZEPS Off
MLS 308/2051 33606 Off
The beta angle reached a minimum angle of 0.0 degrees on 07 November 1997 (DOY
311). The beta angle is now increasing toward a maximum angle of 78.1 degrees
on 27 November 1997 (DOY 331).
Version 62B of the On Board Computer (OBC) flight software was successfully
uplinked on Thursday, 06 November 1997 (310/20:41-20:49 UTC, Orbit 33636). The
flight software revision corrects minor problems identified during the last
yaw maneuver (see AIR 97-113). The revised software functioned flawlessly
during the yaw maneuver.
A Forward-to-Reverse Yaw Around maneuver (number 63) was successfully executed
in Orbit 33646 at 13:27:11 UTC on 07 November 1997. The SSPP Instruments
(ACRIM, SOLSTICE, and SUSIM) were enabled. Due to the SA position at 269
degrees, the effective night length of the Spacecraft after the maneuver was
48 minutes. Post maneuver analysis of data indicate that during the Yaw
Around, the end-of-night load bus voltage was 23.2 volts and the average depth
of discharge (DOD) on both Batteries 2 and 3 was approximately 25.0 percent.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.2 end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 degrees C. Additional battery performance data for each battery on
01 November 1997 and 07 November 1997 are provided in Appendix D.
The clock error ranged from -11.1 to +10.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 15
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 03
% MLS 02
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
TRANSPONDER FREQUENCIES:
Transponder A: 97/301 - 2287.497086
Transponder B: 97/302 - 2287.499098
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 97/310 show the observatory data loss to be 20 hours, 41
minutes, 03.985 seconds (no change). This is a 0.0385 percent data loss which
equals a 99.9615 percent data capture for the mission.
No testing of the new CMS software release (R7B1V2) occurred this week.
Investigation of the stored command verification problem - time to complete -
associated with Applications Processor (AP) software release 10.08 continued.
The cause has been identified, but rectifying the problem continues.
OTHER SIGNIFICANT EVENTS
Completed testing of the new flight software revision (V62B) on Wednesday, 05
November 1997. The flight software revision corrects minor problems identified
during the last yaw maneuver (see AIR 97-113).
Monitoring of the magnetic field model performance continues.
FOT continues to develop an updated Power Profile for UARS due to the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 17 December 1997
at 311/132708.
The Replacement Level Zero Processor (LZPR) will be elevated to operational
status on 15 November 1997. The replacement procedure will result in a planned
10-hour downtime during which no Level Zero Data will be processed - no Near
Realtime Data (NRT) or FOT plots.
The UARS Off-Line personnel will relocate to the second floor of Building 32
adjacent to the UARS/TOMS Mission Operations Room on 10 November 1997.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
CRTs, computers, and other equipment are being installed in the semi-circular
console within the UARS/TOMS Mission Operations Room (MOR). Operations
continue from a dedicated room in the EOS area which is across the hall.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 November 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
33563 305/2223 97-546 HRDI HRDI HTR DELAY TO 450
33564 306/0114 97-547 HRDI HRDI HTR DELAY TO 25
33565 306/0252 97-548 MLS Power ON
33571 306/1220 97-549 HRDI HRDI HTR DELAY TO 450
33606 308/2051 97-550 MLS Power OFF
33617 309/1409 97-551 OBC Clk Adj fr 0.00 to 20.6 msec/day
33634 310/1733 97-552 OBC Clk Adj fr 20.6 to 0.00 msec/day
33636 310/2039 97-553 OBC Load patch for skew wheel reversal
problem
33648 311/1540 97-554 SUSIM Reset sunrise and sunset times
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 November 1997
The following Out-of-Limit (ALARM) conditions were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSVLBEON R-LO Load Bus Volts at EON Continuous
OBSBATVEON2 R-LO Battery 2 voltage at EON Continuous
OBSBATVEON3 R-LO Battery 3 voltage at EON Continuous
OBSVLBEON Y-LO Load Bus Volts at EON 33616-23
OBSBATVEON2 Y-LO Battery 2 voltage at EON 33616-23
OBSBATVEON3 Y-LO Battery 3 voltage at EON 33616-23
OBSSOC2 Y-LO Battery 2 Current SOC 33557,63,78,86,92,
33606,8,17,23,32,8,45,8
MLSCANACTTMP R-LO Scan Actuator temp. 33608,9
WIEUTEMP Y-LO WINDII EU Temperature 33638-53
SOSTMPRNTRN Y-LO Temperature Entrance 33647,8
ACTEMPANA2 Y-LO Temp ACR 2 33647,8
NBRECBI Y-HI NBTR B RECORD CURRENT 33648
PWLBV R-LO CURRENT LBV 33648
PWBAT2V R-LO BATT. 2 CURRENT VOLTAGE 33648
PWBAT3V R-LO BATT. 3 CURRENT VOLTAGE 33648
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 33646
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 November 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
33617 309/14:13 (05 Nov 97) from 00.0 to 20.6 msec/day
33634 310/17:37 (06 Nov 97) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 07 November 1997
01 Nov. 1997 - GMT Day 305 Beta = 22.7 deg, SA Pos.= 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.51 +0.0/ +0.0 18.7 0.000 0.0
2 +78.4/ +0.0 +7.74/+6.44 +31.0/-16.0 24.2 1.022 24.1
3 +67.2/ -5.6 +6.44/+5.15 +30.0/-15.4 24.2 1.023 23.3
07 Nov. 1997 - GMT Day 311 Beta = 3.0 deg, SA Pos.= 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.7 0.000 0.0
2 +44.8/ -22.4 +7.28/+5.79 +30.5/-15.9 23.4 1.019 23.9
3 +50.4/ -11.2 +5.79/+4.51 +29.5/-15.4 23.2 1.020 23.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 November 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-119 FOT 33569 306/0809 14170 Generic Late Acquisition
97-120 FOT 33570 306/0952 N/A RTS 001 tripped
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-111 FOT FIVE 14170 OCT GENERIC LATE ACQ
97-118 FOT 33518 302/23:22 19566 NEG ACQ
97-120 FOT 33570 306/09:52 TMON 01 TRIP-LO VOLT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period